# Imaging a semi-classical horizonless compact object with strong redshift

### v1

@@ I Introduction
In Einstein's theory of general relativity (GR), the gravitational phenomena in the universe are interpreted as the curvature effects of spacetime that are induced by the distribution of matter fields. One direct consequence of such an interpretation is the gravitational lensing effects, i.e., light rays are bent when passing around a massive object, which are important in the contexts of gravitational physics Cunha and Herdeiro (2018), astrophysics Bartelmann (2010); Wambsganss (1998), and cosmology Hoekstra and Jain (2008).
@@
In particular, the gravitational lensing effects around a black hole are so strong that the images of a black hole illuminated by its surrounding light emission are very different from those of usual massive objects. More explicitly, the strong gravitational effects allow the photon trajectories to be extremely curved such that the light rays can orbit the black hole multiple times, or even infinite times, when propagating near it. The infinitely lensed photons correspond to a set of unstable spherical photon orbits around the black hole. The observed black hole image is thus characterized by a bright ring that consists of the intensity of lensed photons, on top of the foreground direct emission Gralla et al. (2019). On the other hand, if the event horizon exists, the photons entering it would never be accessible from outside. Therefore, the bright ring on the image plane would encircle a dark region from which if we trace photon trajectories backward, they would reach the horizon. A bright ring encircling a region with central brightness depression, i.e., the shadow, is an important feature of the images of a black hole in (classical) GR Luminet (1979); Falcke et al. (2000); Bozza (2010); Cunha and Herdeiro (2018). The recent progress in imaging supermassive black holes Akiyama et al. (2019, 2022), achieved by the Event Horizon Telescope collaborations, does confirm such a feature. Furthermore, with the prowess of imaging techniques, we have been ushered into a new era in which probing horizon-scale physics through electromagnetic radiation is possible.
@@
Certainly, the lensing scenario discussed above is for classical black holes. It is well known that GR still has its own issues, e.g. existence of singularities, incompatibility with quantum theories, etc, whose resolution may require quantum gravitational corrections to come into play. In particular, one of the puzzling issues of the quantum nature of black holes is the information paradox Hawking (1976), which is not resolved yet. In this sense, the identity of black holes consistent with quantum theory is still unknown. Essentially, the problem is related to the presence of horizons Mathur (2009); Raju (2022), and therefore, the possibility that large quantum corrections may appear at the horizon scale and even avoid the formation of a horizon Mazur and Mottola (2023); Mathur (2005); Barcelo et al. (2008); Almheiri et al. (2013); Kawai et al. (2013); Barceló et al. (2016), should be seriously considered. It thus becomes crucial to investigate whether one can probe these horizon-scale quantum effects via the shadow images of these quantum-corrected black holes Held et al. (2019); Liu et al. (2020); Brahma et al. (2021); Afrin et al. (2023); Eichhorn et al. (2023a); Zhang et al. (2023).
@@
As we have mentioned, the ring-like feature generated by lensed photons in black hole images is tightly related to the existence of the set of unstable spherical photon orbits around the black hole, rather than the event horizon. Therefore, a horizonless compact object can generate a similar ring-like structure in its image as long as the object is sufficiently compact such that unstable photon orbits exist. The most distinctive feature when there is no event horizon is that some light rays may be reflected at the surface of the compact object, or may propagate through the interior of the object (see Ref. Cardoso and Pani (2019) for a review about testing horizonless compact objects via observations). If these photons pick up additional intensities during their propagation and reach the observer, they can potentially contribute extra rings inside the major bright ring on the image, hence enhancing the central brightness. Such features have been identified in the models of wormholes Ohgami and Sakai (2015); Shaikh et al. (2019a); Paul et al. (2020); Vincent et al. (2021); Wielgus et al. (2020); Wang et al. (2020); Peng et al. (2021); Olmo et al. (2022); Eichhorn et al. (2023b); Guerrero et al. (2022); Delijski et al. (2022); Olmo et al. (2023), boson stars Cunha et al. (2017); Olivares et al. (2020); Rosa and Rubiera-Garcia (2022), gravastars Sakai et al. (2014), fluid stars Rosa (2023), fuzzballs Mayerson and Vercnocke (2023), and naked singularities Shaikh et al. (2019b); Gyulchev et al. (2019); Shaikh and Joshi (2019); Gyulchev et al. (2020). Improving future dynamic range, which, roughly speaking, quantifies the level of difference between the brightest and dimmest image features, may help to distinguish such horizonless compact objects from a black hole covered by a horizon Carballo-Rubio et al. (2022).
@@
However, the modeling of the images of quantum-corrected compact objects without horizons has to be executed with care. In particular, when the photons are allowed to enter the object, some subtle effects that should happen inside the object have to be carefully taken into account. These should include the strong redshifts that can retard the photon propagation as seen by a distant observer, as well as possible interactions between light rays and the internal quantum structure of the object. Intuitively, the combination of these two factors will effectively block the photons propagating inside the object. As a result, the intensity of the inner rings will be suppressed, making the images closer to a classical black hole. Modeling the internal interactions and their associated darkening effects is challenging and is often highly model-dependent due to the lack of knowledge about how to describe such quantum interactions based on microscopic dynamics. For example, an effective approach of darkening images motivated by interactions was recently proposed in the construction of the images of fuzzballs Bacchini et al. (2021) and compact topological solitons Heidmann et al. (2023). It was shown that the shadow images cast by these horizonless compact objects can be almost indistinguishable from that of a classical black hole. Gravitational redshifts, on the other hand, should be able to darken the images in a more model-independent manner. This is because gravity is determined by a given energy-momentum distribution, independent of the details of the microscopic components. Therefore, we can expect that an extremely strong redshift, if any, should darken the images of a horizonless compact object in a way that does not depend much on the details of its internal interactions. In this paper, we would like to examine this possibility and study how generic such a darkening process is and how much the darkened images mimic the images of a classical black hole.
@@
This is the self-consistent equation in a mean-field approximation of quantum gravity, where gravity is described in terms of a classical metric $g_{\mu\nu}$, and matter is represented by quantum operators Birrell and Davies (1984); Kiefer (2012). In this framework, a macroscopic object is described as a collection of many excited quanta in an excited state $|\psi\rangle$, whose energy-momentum $\langle\psi|T_{\mu\nu}|\psi\rangle$ produces the self-gravity $g_{\mu\nu}$ according to Eq. (1). In Ref. Kawai et al. (2013); Kawai and Yokokura (2015, 2016); Ho (2017); Kawai and Yokokura (2017, 2020, 2022), by solving Eq. (1) self-consistently, a 4D spherically symmetric spacetime region was obtained as the configuration formed in the collapse of a spherical matter, where the backreaction of particles created during the collapse was considered. It represents a compact object with an outer surface $R_{\textrm{out}}$ located just outside the Schwarzschild radius $a_{0}(\gg l_{p}\equiv\sqrt{\hbar G})$ instead of a horizon, and it evaporates due to Hawking-like radiation in the timescale $\Delta t\sim a_{0}^{3}/l_{p}^{2}$. For a generic collapse, a dense structure with an exponentially large redshift and near-Planckian curvatures is formed around the surface, while the structure in deeper regions depends on the details of the initial distribution of the collapsing matter Kawai and Yokokura (2016). For an adiabatic formation in a heat bath at Hawking temperature, the dense structure continues inward (except for a small center part) to form a radially uniform dense configuration, which can be considered the most thermodynamically typical Kawai and Yokokura (2015, 2022). In this paper, we focus on the uniform dense configuration and dub it the quantum horizonless compact object (QHCO). In addition, one can evaluate the thermodynamic entropy of the interior by a thermodynamic method Kawai and Yokokura (2016); Yokokura (2022, 2023) and a statistical-mechanical one Kawai and Yokokura (2020), to reproduce the Bekenstein-Hawking formula Bekenstein (1973); Hawking (1975). [^1] Here, the strong self-gravity plays an essential role in changing the entropy from the volume law to the area law Yokokura (2022). It was later found Yokokura (2023) that the QHCO is the entropy-maximized configuration that saturates the entropy bounds proposed by Bekenstein Bekenstein (1981) and Bousso Bousso (1999). Thus, QHCO has (some of) properties that a black hole should have in quantum theory and could be a candidate for a quantum black hole.
@@
We would like to emphasize again that the QHCO model satisfies self-consistently the semi-classical Einstein equations (1) with many matter fields. [^2] Indeed, one can use a technique based on the properties of the interior metric, evaluate the renormalized energy-momentum tensor $\langle\psi|T_{\mu\nu}|\psi\rangle$ for, say, many massless scalar fields, and equate it to $G_{\mu\nu}$, to solve Eq. (1) self-consistently Kawai and Yokokura (2020). As a result, the relation (7) is obtained. The high curvature inside induces quantum fluctuations of quanta, generating a large tangential pressure $\langle T^{\theta}{}_{\theta}\rangle$ Kawai and Yokokura (2020) that supports the configuration against the strong self-gravity. [^3] This is precisely why QHCOs can evade the Buchdahl limit Buchdahl (1959) and reach extreme compactness, similar to the case of very anisotropic stars Raposo et al. (2019). Although the model obeys the semi-classical description, it non-trivially represents a non-perturbative solution in $\hbar$ in the sense that the curvatures as well as the metric itself cannot exist in the limit $\hbar\rightarrow 0$. Indeed, it belongs to a branch in the solution space of Eq. (1) different from those that include classical black hole metrics Kawai and Yokokura (2022); Ho et al. (2023).
@@
The main theme of this paper is to investigate the images of QHCOs. Here, we neglect the time evolution of the evaporation and consider them static. When illuminated by light emission, how alike, or how different, are the shadow images cast by QHCOs compared with those of a classical black hole? As we will show later, light rays are allowed to propagate through the interior of QHCOs due to the absence of the event horizon. If these light rays can reach the observer, they can generate additional bright rings inside the major ring on the images. However, in reality, these light rays will take an extremely long time during their passage in the dense configuration due to the exponentially strong redshift. The delays of these light rays then effectively darken the observed images, such that the darkened images become almost indistinguishable from those of classical black holes. We expect this feature to be insensitive to the details of possible internal interactions. In fact, the QHCO metric incorporates the internal interaction effects through a phenomenological parameter $\eta$ (see sec. II), and we show that the image becomes closer to that of classical black holes for a larger $\eta$, i.e., stronger interactions. These render the QHCO a perfect model for (classical) black hole mimickers. In addition, we find that a tiny feature of the QHCO image, which is more pronounced for larger $n$, will appear if the light source exists just inside the QHCO surface. This prediction is independent of the details of the emission profile, as long as the inner part of the surface is lit up. Therefore, such a feature may be a test-bed for near-horizon quantum effects with future improvements in observational techniques, e.g., substantially high dynamic range.
This paper is organized as follows. In sec. II, we review the QHCO spacetime in a more explicit manner, including the metric and some of the properties of QHCOs. In sec. III, we investigate in detail the photon geodesics of QHCO spacetimes. Considering a QHCO surrounded by an optically and geometrically thin accretion disk, we investigate the ``ideal" QHCO images in sec. IV based on the assumption that the light rays that propagate through QHCOs can reach the observer, no matter how long they may take inside QHCOs as seen by the observer. Then, in sec. V, we generate a set of more realistic QHCO images by including the darkening effects caused by the strong redshifts inside QHCOs. We discuss how the images may differ from, or mimic, the images of a classical Schwarzschild black hole. Finally, in sec. VI, we conclude the paper by summarizing the results obtained and discussing the implications of the $(\eta,n)$-dependence of the images to fundamental aspects of quantum gravity.
@@ II Quantum horizonless compact object
More explicitly, the QHCO spacetime can be described by the following spherically symmetric and static metric Kawai et al. (2013); Kawai and Yokokura (2015, 2016, 2017, 2020, 2022); Yokokura (2022, 2023):
@@
Now, we fix the constant $C$ by using Israel's junction condition Poisson (2009), that is, the continuity of the metric functions $A_{1}(r_{1})$ and $A_{2}(r_{2})$ at the surface $R_{\textrm{out}}$:
@@ IV Ideal QHCO images
As discussed above, the absence of the event horizon for the QHCO model allows the existence of photons that have traveled through QHCOs before being observed. It is then interesting to understand how these photons may affect, in principle, the images of QHCOs as seen by a distant observer. In this section, we study a very idealized scenario in which such penetrating photons reach the observer no matter how strongly they are redshifted inside QHCOs. Then, we show such ``ideal" images of QHCOs, which have different image features from those of classical Schwarzschild black holes, although they are not physical in the sense that the effect of a physical timescale is not considered here (see sec. V).
@@
Let us construct the formula of the observed intensity $I_{o}$ based on the above assumptions. First, because there is no absorption, the specific intensities $I_{\nu_{e}}$ and $I_{\nu_{o}}$ with frequencies $\nu_{e}$ and $\nu_{o}$ in the emission frame and observer's reference frame, are related generically via Lindquist (1966)
@@
Now, we are ready to plot the observed intensity $I_{o}$ and construct the corresponding ideal QHCO images. We first solve the equations of motion (17) and (27) to obtain the trajectories, from which we identify $r_{k}$ for each of them; we apply the intensity formula (IV) with the GLM model (35) and obtain the intensity $I_{o}(b)$; and then we ``rotate" it around the origin because of the spherical symmetry to get the images. The results are given by Fig. 7 and Fig. 8 for the GLM1 and GLM2 emission profiles, respectively. [^11] The intensity profiles and the images of the GLM3 emission are qualitatively very similar to those of GLM1. Therefore, we do not show them here. Here, the top and lower panels represent, respectively, the observed intensities $I_{o}(b)$ and the images for the classical black hole and the QHCOs.

### v2

@@ I Introduction
One prediction in General Relativity is the gravitational lensing effects, i.e., light rays are bent when passing around a massive object, which are important in the contexts of gravitational physics Cunha and Herdeiro (2018), astrophysics Bartelmann (2010); Wambsganss (1998), and cosmology Hoekstra and Jain (2008). In particular, the lensing effects around a black hole are so strong that the images of a black hole illuminated by its surrounding light emission are very different from those of usual massive objects. More explicitly, the photon trajectories can be extremely curved such that the light rays can orbit the black hole multiple times, or even infinite times, when propagating near it. The infinitely lensed photons correspond to a set of unstable spherical photon orbits around the black hole. The observed black hole image is thus characterized by a bright ring that consists of the intensity of lensed photons, on top of the foreground direct emission Gralla et al. (2019). On the other hand, the photons entering the event horizon cannot be accessed from outside. Therefore, the bright ring on the image plane would encircle a dark region that corresponds to photon trajectories terminating at the horizon. A bright ring encircling a region with central brightness depression, i.e., the shadow, is an important feature of the images of a black hole in (classical) GR Luminet (1979); Falcke et al. (2000); Bozza (2010); Cunha and Herdeiro (2018). The recent progress in imaging supermassive black holes Akiyama et al. (2019, 2022) achieved by the Event Horizon Telescope collaborations has confirmed such a feature.
@@
The lensing scenario discussed above is for classical black holes. It is well known that GR still has its own issues, e.g. existence of singularities, incompatibility with quantum theories, etc, whose resolution may require quantum gravitational corrections. In particular, one unresolved puzzle of the quantum nature of black holes is the information paradox Hawking (1976), which indicates that the identity of black holes consistent with quantum theory is still unknown. Essentially, the problem is related to the presence of horizons Mathur (2009); Raju (2022), and therefore, the possibility that large quantum corrections may appear at the horizon scale and even avoid the formation of a horizon Mazur and Mottola (2023); Mathur (2005); Barcelo et al. (2008); Almheiri et al. (2013); Kawai et al. (2013); Barceló et al. (2016), should be seriously considered. It thus becomes crucial to investigate whether one can probe these horizon-scale quantum effects via the shadow images of these quantum-corrected black holes Held et al. (2019); Liu et al. (2020); Brahma et al. (2021); Afrin et al. (2023); Eichhorn et al. (2023a); Zhang et al. (2023).
@@
Theoretically, a horizonless compact object can also generate a ring-like structure in its image as long as it is sufficiently compact such that unstable photon orbits exist. The most distinctive feature when there is no event horizon is that some light rays may be reflected at the surface of the compact object, or may propagate through its interior (see Ref. Cardoso and Pani (2019) for a review about testing horizonless compact objects via observations). If these photons pick up additional intensities during their propagation and reach the observer, they may contribute extra rings inside the major bright ring on the image, hence enhancing the central brightness. Such features have been identified in the models of wormholes Ohgami and Sakai (2015); Shaikh et al. (2019a); Paul et al. (2020); Vincent et al. (2021); Wielgus et al. (2020); Wang et al. (2020); Tsukamoto (2021a); Peng et al. (2021); Olmo et al. (2022); Eichhorn et al. (2023b); Guerrero et al. (2022); Delijski et al. (2022); Olmo et al. (2023), boson stars Cunha et al. (2017); Olivares et al. (2020); Rosa and Rubiera-Garcia (2022), gravastars Sakai et al. (2014), fluid stars Rosa (2023), fuzzballs Mayerson and Vercnocke (2023), and naked singularities Shaikh et al. (2019b); Gyulchev et al. (2019); Shaikh and Joshi (2019); Gyulchev et al. (2020); Stashko et al. (2021); Stashko and Zhdanov (2021); Tsukamoto (2021b); Huang et al. (2024). Improving future dynamic range, which, roughly speaking, quantifies the level of difference between the brightest and dimmest image features, may help to distinguish such horizonless compact objects from a black hole covered by a horizon Carballo-Rubio et al. (2022).
@@
However, when the photons are allowed to enter the object, some subtle effects inside have to be carefully considered. These include the strong redshifts that can retard the photon propagation as seen by a distant observer, as well as possible interactions between light rays and the internal quantum structure of the object. Intuitively, the combination of these two factors will effectively block the photons propagating inside the object, hence suppressing the intensity of the inner rings, and resembling more the classical black hole images. Modeling the internal interactions and their associated darkening effects is challenging and is often model-dependent. Recently, an effective approach to darkening images was proposed in the construction of the images of fuzzballs Bacchini et al. (2021) and compact topological solitons Heidmann et al. (2023). It was shown that the shadow images of these horizonless compact objects can be almost indistinguishable from classical black hole images. Gravitational redshifts, on the other hand, can darken images in a more model-independent manner because gravity is determined by a given energy-momentum distribution, independent of the details of the microscopic components. Therefore, the presence of extremely strong redshift should darken the images of a horizonless compact object in a way that does not depend much on the details of its internal interactions. In this paper, we will examine how generic such a darkening process is and how much the darkened images mimic the images of a classical black hole.
@@
This is the self-consistent equation in a mean-field approximation of quantum gravity, where gravity is described by a classical metric $g_{\mu\nu}$, and matter is represented by quantum operators Birrell and Davies (1984); Kiefer (2012). In this framework, a macroscopic object is described as a collection of many excited quanta in an excited state $|\psi\rangle$. In Ref. Kawai et al. (2013); Kawai and Yokokura (2015, 2016); Ho (2017); Kawai and Yokokura (2017, 2020, 2022), by solving Eq. (1) self-consistently, a 4D spherically symmetric spacetime region was obtained as the configuration formed in the collapse of a spherical matter, where the backreaction of particles created during the collapse was considered. It represents a compact object with an outer surface $R_{\textrm{out}}$ located just outside the Schwarzschild radius $a_{0}(\gg l_{p}\equiv\sqrt{\hbar G})$ instead of a horizon, and it evaporates due to Hawking-like radiation in the timescale $\Delta t\sim a_{0}^{3}/l_{p}^{2}$. For a generic collapse, a dense structure with an exponentially large redshift and near-Planckian curvatures is formed around the surface, while the structure in deeper regions depends on the details of the initial distribution of the collapsing matter Kawai and Yokokura (2016). For an adiabatic formation in a heat bath at Hawking temperature, the dense structure continues inward (except for a small center part) to form a radially uniform dense configuration, which can be considered the most thermodynamically typical Kawai and Yokokura (2015, 2022). In this paper, we focus on the uniform dense configuration and dub it the quantum horizonless compact object (QHCO). In addition, one can evaluate the thermodynamic entropy of the interior by a thermodynamic method Kawai and Yokokura (2016); Yokokura (2024) and a statistical-mechanical one Kawai and Yokokura (2020), to reproduce the Bekenstein-Hawking formula Bekenstein (1973); Hawking (1975). [^1] Here, the strong self-gravity plays an essential role in changing the entropy from the volume law to the area law Yokokura (2024). It was later found Yokokura (2023) that the QHCO is the entropy-maximized configuration that saturates the entropy bounds proposed by Bekenstein Bekenstein (1981) and Bousso Bousso (1999). Thus, QHCO has some properties that a black hole should have in quantum theory and could be a candidate for a quantum black hole.
@@
We would like to emphasize again that the QHCO model satisfies self-consistently the semi-classical Einstein equations (1) with many matter fields. [^2] Indeed, one can use a technique based on the properties of the interior metric, evaluate the renormalized energy-momentum tensor $\langle\psi|T_{\mu\nu}|\psi\rangle$ for, say, many massless scalar fields, and equate it to $G_{\mu\nu}$, to solve Eq. (1) self-consistently Kawai and Yokokura (2020). As a result, the relation (7) is obtained. The high curvature inside induces quantum fluctuations of quanta, generating a large tangential pressure $\langle T^{\theta}{}_{\theta}\rangle$ Kawai and Yokokura (2020) that supports the configuration against the strong self-gravity. [^3] This is precisely why QHCOs can evade the Buchdahl limit Buchdahl (1959) and reach extreme compactness, similar to the case of very anisotropic stars Raposo et al. (2019). Although the model obeys the semi-classical description, it non-trivially represents a non-perturbative solution in $\hbar$ in the sense that the curvatures and the metric itself cannot exist in the limit $\hbar\rightarrow 0$. Indeed, it belongs to a branch in the solution space of Eq. (1) different from those that include classical black hole metrics Kawai and Yokokura (2022); Ho et al. (2023).
@@
The main theme of this paper is to investigate the images of QHCOs neglecting the time evolution of the evaporation. As we will show later, light rays are allowed to propagate through the interior of QHCOs due to the absence of the event horizon. If these light rays can reach the observer, they can generate additional bright rings inside the major ring on the images. However, in reality, these light rays will take an extremely long time during their passage in the dense configuration due to the exponentially strong redshift. The delays of these light rays then effectively darken the observed images, such that the darkened images become almost indistinguishable from those of classical black holes. We expect this feature to be insensitive to the details of possible internal interactions. In fact, the QHCO metric incorporates the internal interaction effects through a phenomenological parameter $\eta$ (see sec. II), and we show that the image becomes closer to that of classical black holes for stronger interactions. These render the QHCO a perfect model for black hole mimickers. In addition, we find that a tiny feature of the QHCO image, which is more pronounced for larger $n$, will appear if light sources exist just inside the QHCO surface. This prediction is independent of the details of the emission profile, as long as the inner part of the surface is lit up. Therefore, such a feature may be a test-bed for near-horizon quantum effects with future improvements in observational techniques, e.g., substantially high dynamic range.
This paper is organized as follows. In sec. II, we review the QHCO spacetime more explicitly, including the metric and some properties of QHCOs. In sec. III, we investigate the photon geodesics of QHCO spacetimes. Considering a QHCO surrounded by an optically and geometrically thin accretion disk, we investigate the “ideal” QHCO images in sec. IV assuming that the light rays propagating through QHCOs can reach the observer, no matter how long they may take inside QHCOs as seen by the observer. Then, in sec. V, we generate a set of more realistic QHCO images by including the darkening effects caused by the strong redshifts inside QHCOs. We discuss how the images may differ from, or mimic, the images of a classical Schwarzschild black hole. Finally, in sec. VI, we conclude the paper by summarizing the results obtained and discussing their implications to fundamental aspects of quantum gravity.
@@ II Quantum horizonless compact object
More explicitly, the QHCO spacetime can be described by the following spherically symmetric and static metric Kawai et al. (2013); Kawai and Yokokura (2015, 2016, 2017, 2020, 2022); Yokokura (2024, 2023):
@@
Now, we fix the constant $C$ by using Israel’s junction condition Poisson (2009), that is, the continuity of the metric functions $A_{1}(r_{1})$ and $A_{2}(r_{2})$ at the surface $R_{\textrm{out}}$:
@@ IV Ideal QHCO images
As discussed above, the absence of the event horizon for the QHCO model allows the existence of photons that have traveled through QHCOs before being observed. It is then interesting to understand how these photons may affect, in principle, the images of QHCOs as seen by a distant observer. In this section, we study a very idealized scenario in which such penetrating photons reach the observer no matter how strongly they are redshifted inside QHCOs. Then, we show such “ideal” images of QHCOs, which have different image features from those of classical Schwarzschild black holes, although they are not physical in the sense that the effect of a physical timescale is not considered here (see sec. V).
@@
Let us construct the formula of the observed intensity $I_{o}$ based on the above assumptions. First, because there is no absorption, the specific intensities $I_{\nu_{e}}$ and $I_{\nu_{o}}$ with frequencies $\nu_{e}$ and $\nu_{o}$ in the emission frame and observer’s reference frame, are related generically via Lindquist (1966)
@@
Now, we are ready to plot the observed intensity $I_{o}$ and construct the corresponding ideal QHCO images. We first solve the equations of motion (17) and (27) to obtain the trajectories, from which we identify $r_{k}$ for each of them; we apply the intensity formula (IV) with the GLM model (35) and obtain the intensity $I_{o}(b)$; and then we “rotate” it around the origin because of the spherical symmetry to get the images. The results are given by Fig. 7 and Fig. 8 for the GLM1 and GLM2 emission profiles, respectively. [^11] The intensity profiles and the images of the GLM3 emission are qualitatively very similar to those of GLM1. Therefore, we do not show them here. Here, the top and lower panels represent, respectively, the observed intensities $I_{o}(b)$ and the images for the classical black hole and the QHCOs.
Refer to captionRefer to captionRefer to caption